# MoodAngels: A Retrieval-augmented Multi-agent Framework for Psychiatry Diagnosis

### v1

@@ Abstract
The application of AI in psychiatric diagnosis faces significant challenges, including the subjective nature of mental health assessments, symptom overlap across disorders, and privacy constraints limiting data availability. To address these issues, we present MoodAngels, the first specialized multi-agent framework for mood disorder diagnosis. Our approach combines granular-scale analysis of clinical assessments with a structured verification process, enabling more accurate interpretation of complex psychiatric data. Complementing this framework, we introduce MoodSyn, an open-source dataset of 1,173 synthetic psychiatric cases that preserves clinical validity while ensuring patient privacy. Experimental results demonstrate that MoodAngels outperforms conventional methods, with our baseline agent achieving 12.3% higher accuracy than GPT-4o on real-world cases, and our full multi-agent system delivering further improvements. Evaluation in the MoodSyn dataset demonstrates exceptional fidelity, accurately reproducing both the core statistical patterns and complex relationships present in the original data while maintaining strong utility for machine learning applications. Together, these contributions provide both an advanced diagnostic tool and a critical research resource for computational psychiatry, bridging important gaps in AI-assisted mental health assessment [^1] Code and synthetic data sample are available in https://github.com/elsa66666/MoodAngels..
@@ 1 Introduction
Beyond diagnostic frameworks, clinical data scarcity presents another research challenge. Despite longstanding interest in depression and bipolar detection, existing methods predominantly analyze social media posts [16, 17, 18], which may be distorted by social norms or exaggeration. To enable accurate AI-driven psychiatric diagnosis and early detection, we construct the open-source synthetic dataset MoodSyn with 1,173 synthetic psychiatric cases, containing: selected five groups of top-related scale data, total scores from 13 common mental disorder scales, and mood disorder labels. Through a comprehensive evaluation of quality, ML efficiency, and privacy protection, we demonstrate that this synthetic dataset maintains high fidelity while safeguarding visitor confidentiality.
@@ 2 Retrieval-augmented Multi-agent Framework
In this section, we propose a retrieval-augmented multi-agent framework tailored for the diagnosis of mood disorders, integrating structured clinical knowledge with dynamic decision-making. The framework consists of three specialized diagnostic agents, each performing an independent diagnostic process based on reorganized medical records and selected diagnostic scales. These agents employ two core analytical tools: a Granular-Scale Analysis module that matches symptoms to diagnostic criteria at a fine-grained level, and a Case-Based Retrieval module that retrieves relevant prior cases from a curated knowledge base. The diagnostic opinions of the agents are then synthesized through a structured debate mechanism to form a final judgment.
@@ 2.1 Granular-scale Analysis
To address this challenge, we introduce an innovative granular-scale analysis approach. Rather than evaluating scales solely through total scores, we decompose them into individual item-level responses. We employed a comprehensive set of clinical scales to assess various aspects of mental health, including eight self-reported scales [^2] The self-reported scales are available at: CTQ, DAS, GAD-7, HCL-32, MDQ, NSSI, PHQ-9 and SHAPS. and five clinician-evaluated scales [^3] The clinician-administered scales are available at: BPRS, HAMA, HAMD-24, MCCB and YMRS. In this study, we employ the Chinese version of the HAMD-24, which features a different question order compared to the English version.. These tools provide valuable insights into both subjective experiences and symptoms, as well as clinical observations and structured evaluations of clients’ mental states and behaviors. All scales are routinely used in hospital settings for clinical diagnosis, ensuring the reliability and validity of the collected data.
To identify the most relevant questions for mood disorder diagnosis, we computed the Pearson correlation between each question’s score (and total score) and the presence of a mood disorder, selecting the top 5% with the highest correlations. These questions naturally clustered into key symptom groups: depressive mood, loss of interest, anxiety, insomnia, and suicidal tendencies. These groups enhance diagnostic robustness through cross-validation and comprehensive symptom coverage. To further refine our framework, we included clinically significant PHQ-9 questions, such as phq9_Q2 (depressed mood) and phq9_Q1 (loss of interest), even if their correlation scores were slightly below the threshold, ensuring a nuanced and reliable diagnostic process (see Appendix D).
@@ 2.2 Retrieval Datastore
Since overlapping symptoms may correspond to multiple disorders, we extract and structure diagnostic and differential criteria from the Diagnostic and Statistical Manual of Mental Disorders: DSM-5 [15], a widely recognized authority in psychiatry, to build a retrievable knowledge base. The knowledge base construction process is detailed in Appendix C.
To prevent MoodAngels from making arbitrary decisions based solely on symptom presentation, we also incorporate clinicians’ diagnostic expertise by including anonymized clinical data for retrieval. These experiences are also beneficial when a visitor’s symptoms are ambiguous, for historical diagnostic precedents may offer additional interpretive insights. The clinical data used in our study consists of anonymized real-world hospital cases, totaling 2804 entries. We partitioned the dataset such that 80% of the cases are used as historical cases for retrieval, while the remaining 20% serve as the test set. All clients in the dataset have completed scale assessments, although clients without diagnosed conditions do not have medical records available, and our agents are not pre-informed about this distinction. The dataset statistics are summarized in Table 1. Appendix E details the processing steps applied to both medical records and scale data, which aim to optimize the performance of LLM-based agents for more accurate analysis.
@@ 2.3 Diagnostic Agents
To mitigate overreliance on past cases (which could overlook individual variability in psychiatric diagnoses), we develop three diagnostic variants with differing levels of historical dependence: Angel.R (no reference to previous cases), Angel.D (displays retrieved cases as context), and Angel.C (compares each retrieved case with the current query and returns an analysis as context).
By aggregating independent diagnoses from these three agents and facilitating debate among their conclusions, our final diagnosis model, multi-Angels, bridges the gap between computational decision-making and the nuanced understanding essential for accurate psychiatric evaluations.
The following parts introduce the main components of our agents:
Symptom Matching To align client symptoms in medical records with DSM-5 diagnostic criteria, we process records and compute relevance between records and criteria using dense vector encoding. The BGE-M3 embedder [19] is employed for its strong semantic embedding capabilities. We retrieve the top-5 most similar criteria, returning their text, classification, and similarity scores (detailed in Appendix G.1). The tool does not diagnose but provides results for agent analysis, ensuring decisions integrate quantitative data and clinical expertise, mitigating over-reliance on single metrics.
@@
Scale Performance Analysis We diagnose the presence of mood disorder using 16 key questions selected in Section 2.1 as the most mood-relevant items. Client performances are converted from numeric scores to textual descriptions based on question content and options. For agent interpretability, performances are reorganized into coherent descriptive paragraphs, enhancing analysis effectiveness.
Similar Cases Retrieval To leverage clinical experience from similar cases, we develop two optional tools for retrieving medical records and scales with similar performance. After performing similarity retrieval, our tools generate different outputs tailored to the type of diagnosis agent in use. For Angel.R, this tool is intentionally excluded to minimize potential interference from the diagnostic outcomes of other cases. For Angel.D, the tool directly returns the retrieved cases for reference, enabling the agent to review and draw insights from them. For Angel.C, the tool conducts a detailed comparison of similarities and differences among the retrieved cases and returns an analysis text summarizing the findings. Consistent with the symptom matching tool, we employ BGE-M3 as the retriever and select the top 5 most relevant cases as the retrieved records. This approach ensures that the tool adapts to the specific needs of each diagnosis agent, enhancing the diagnostic process while maintaining flexibility and precision.
Multi-agent Diagnosis To integrate insights from all Angels and improve diagnostics, Angel.R, Angel.D, and Angel.C first provide independent decisions and reasoning. A Judge Agent consolidates their inputs. If consensus is reached, the Judge outputs the diagnosis and reasoning.
For disagreements, two Debate Agents are introduced: a Positive Agent, supporting a mood disorder diagnosis, and a Negative Agent, opposing it. Both Debate Agents and the Judge access symptom matching results, scale performances, relevant cases, and the Angels’ diagnosis and reasoning. In each debate round, the Positive Agent speaks first, followed by the Negative Agent. After each round, the Judge evaluates the arguments and decides whether to conclude the debate. If concluded, the Judge delivers the final diagnosis and supporting reasons.
@@ 3 MoodSyn Dataset
The MoodSyn dataset addresses the critical need for clinically valid yet privacy-preserving data in computational psychiatry by introducing an open-source collection of 1,173 synthetic cases, as the statistics shown in Table 2. Each case captures the essential characteristics of psychiatric assessments through 25 carefully selected features, including 16 diagnostic questions, 8 standard scale scores, and expert-verified mood disorder labels. A data example is shown in Appendix F.1.
@@
Dataset Construction. The MoodSyn dataset is constructed on an advanced synthesis pipeline built upon the TabSyn framework [20], which integrates variational autoencoders with diffusion models through several technical innovations. The architecture employs adaptive tokenization and hierarchical encoding to maintain the complex relationships between clinical features, while a conditional denoising process preserves the nuanced symptom patterns characteristic of mood disorders. During generation, dynamic loss weighting ensures the mathematical consistency between individual question responses and their corresponding scale scores, and context-aware sampling maintains clinically meaningful feature correlations. The construction process is detailed in Appendix F.1.
Dataset Evaluation. Following Zhang et al. [20], we conduct a comprehensive evaluation of MoodSyn across five critical dimensions: statistical density, data quality, machine learning efficacy, privacy preservation, and logistic detectability (see Appendix F.2 for full methodology). The results demonstrate exceptional fidelity to real clinical data, with MoodSyn accurately reproducing both core statistical patterns and complex symptom relationships while maintaining strong utility for machine learning applications. Specifically, our evaluation reveals high-density preservation of univariate and multivariate distributions matching the original data’s feature patterns, robust performance in downstream diagnostic prediction tasks comparable to real data, and near-indistinguishability in rigorous detection tests. Furthermore, MoodSyn provides stronger privacy guarantees than traditional anonymization approaches through its synthetic generation process. This unique combination of statistical fidelity, clinical validity, and privacy protection establishes MoodSyn as both a reliable analytical surrogate and a valuable resource for advancing AI applications in mental health research.
@@ 4.1 Experimental Settings
Baselines. To evaluate the effectiveness of our agent, we compare it against four baseline LLMs: LLaMA3-8B-Instruct [21], Mistral-7B-Instruct-v0.3 [22], GPT-4o (gpt-4o-2024-08-06) [23], DeepSeek-V3 [24]. Each LLM is provided with a standardized input consisting of: (1) A combined medical record, presented as a unified narrative. (2) A summary of scale test results, listing the client’s scores across multiple psychological scales along with their clinical implications. The LLMs are prompted to determine whether the client has a mood disorder (including depression or bipolar disorder) and to provide a structured explanation in JSON format. An example of the prompt used for baseline models is shown in Figure 8.
Hyperparameters. For each model, we employ default parameter settings, utilizing official models for open-source LLMs obtained from Hugging Face or the API from the official website. These testing procedures take place on a computational infrastructure consisting of two NVIDIA A800 Tensor Core GPUs, equipped with 80GB of memory.
@@ 4.2 Analysis of Experimental Results
The experimental results presented in Tables 3 and 4 demonstrate the clear advantages of this approach. All MoodAngels variants show substantial improvements over baseline LLMs, with even our foundational agent achieving 12.3% greater accuracy than GPT-4o, underscoring the effectiveness of our redesigned diagnostic process. A deeper examination of the agent comparisons reveals important insights about psychiatric diagnosis. The performance advantage of Angel.D over Angel.R confirms the clinical value of historical cases as reference points, while the slight dip observed with Angel.C serves as a caution against over-reliance on past cases. Particularly telling are those instances where only one agent succeeds in making the correct diagnosis while others fail, highlighting both the complexity of psychiatric assessment and the complementary nature of different diagnostic approaches.
These findings collectively validate our multi-agent framework’s design, which synthesizes granular symptom analysis, structured diagnostic protocols, and balanced clinical experience integration through its collaborative debate mechanism. The framework’s ultimate performance superiority emerges from this sophisticated combination of innovations, each addressing specific limitations in conventional psychiatric assessment methods while working in concert to achieve more accurate and reliable diagnoses.
@@
Model Sensitivity Accuracy MCC Macro F1
@@
[…]
@@
[…]
@@
Method Recall Accuracy MCC Macro F1
@@
[…]

### v2

@@ Abstract
The application of AI in psychiatric diagnosis faces significant challenges, including the subjective nature of mental health assessments, symptom overlap across disorders, and privacy constraints limiting data availability. To address these issues, we present MoodAngels, the first specialized multi-agent framework for mood disorder diagnosis. Our approach combines granular-scale analysis of clinical assessments with a structured verification process, enabling more accurate interpretation of complex psychiatric data. Complementing this framework, we introduce MoodSyn, an open-source dataset of 1,173 synthetic psychiatric cases that preserves clinical validity while ensuring patient privacy. Experimental results demonstrate that MoodAngels outperforms conventional methods, with our baseline agent achieving 12.3% higher accuracy than GPT-4o on real-world cases, and our full multi-agent system delivering further improvements. Evaluation in the MoodSyn dataset demonstrates exceptional fidelity, accurately reproducing both the core statistical patterns and complex relationships present in the original data while maintaining strong utility for machine learning applications. Together, these contributions provide both an advanced diagnostic tool and a critical research resource for computational psychiatry, bridging important gaps in AI-assisted mental health assessment.
@@ 1 Introduction
Beyond diagnostic frameworks, clinical data scarcity presents another research challenge. Despite longstanding interest in depression and bipolar detection, existing methods predominantly analyze social media posts [16, 17, 18], which may be distorted by social norms or exaggeration. To enable accurate AI-driven psychiatric diagnosis and early detection, we construct the open-source synthetic dataset MoodSyn [^1] Code and synthetic data sample are available in MoodAngels. with 1,173 synthetic psychiatric cases, containing: selected five groups of top-related scale data, total scores from 13 common mental disorder scales, and mood disorder labels. Through a comprehensive evaluation of quality, ML efficiency, and privacy protection, we demonstrate that this synthetic dataset maintains high fidelity while safeguarding visitor confidentiality.
@@ 2 Retrieval-augmented Multi-agent Framework
In this section, we propose a retrieval-augmented multi-agent framework tailored for the diagnosis of mood disorders [^2] The problem definition is provided in Appendix B.1., integrating structured clinical knowledge with dynamic decision-making. The framework consists of three specialized diagnostic agents, each performing an independent diagnostic process based on reorganized medical records and selected diagnostic scales. These agents employ two core analytical tools: a Granular-Scale Analysis module that matches symptoms to diagnostic criteria at a fine-grained level, and a Case-Based Retrieval module that retrieves relevant prior cases from a curated knowledge base. The diagnostic opinions of the agents are then synthesized through a structured debate mechanism to form a final judgment.
@@ 2.1 Granular-scale Analysis
To address this challenge, we introduce an innovative granular-scale analysis approach [^3] We provide simple descriptions of scales and the selection process of relevant questions in Appendix B.2.. Rather than evaluating scales solely through total scores, we decompose them into individual item-level responses. We employed a comprehensive set of clinical scales to assess various aspects of mental health, including eight self-reported scales and five clinician-evaluated scales. These tools provide valuable insights into both subjective experiences and symptoms, as well as clinical observations and structured evaluations of clients’ mental states and behaviors. All scales are routinely used in hospital settings for clinical diagnosis, ensuring the reliability and validity of the collected data.
To identify the most relevant questions for mood disorder diagnosis, we computed the Pearson correlation between each question’s score (and total score) and the presence of a mood disorder, selecting the top 5% with the highest correlations. These questions naturally clustered into key symptom groups: depressive mood, loss of interest, anxiety, insomnia, and suicidal tendencies. These groups enhance diagnostic robustness through cross-validation and comprehensive symptom coverage. To further refine our framework, we included clinically significant PHQ-9 questions, such as phq9_Q2 (depressed mood) and phq9_Q1 (loss of interest), even if their correlation scores were slightly below the threshold, ensuring a nuanced and reliable diagnostic process.
@@ 2.2 Retrieval Datastore
Since overlapping symptoms may correspond to multiple disorders, we extract and structure diagnostic and differential criteria from the Diagnostic and Statistical Manual of Mental Disorders: DSM-5 [15], a widely recognized authority in psychiatry, to build a retrievable knowledge base. The knowledge base construction process is detailed in Appendix B.3.2.
To prevent MoodAngels from making arbitrary decisions based solely on symptom presentation, we also incorporate clinicians’ diagnostic expertise by including anonymized clinical data for retrieval. These experiences are also beneficial when a visitor’s symptoms are ambiguous, for historical diagnostic precedents may offer additional interpretive insights. The clinical data used in our study consists of anonymized real-world hospital cases, totaling 2804 entries. We partitioned the dataset such that 80% of the cases are used as historical cases for retrieval, while the remaining 20% serve as the test set. All clients in the dataset have completed scale assessments, although clients without diagnosed conditions do not have medical records available, and our agents are not pre-informed about this distinction. The dataset statistics are summarized in Table 1.
@@ 2.3 Diagnostic Agents
To mitigate overreliance on past cases (which could overlook individual variability in psychiatric diagnoses), we develop three diagnostic variants with differing levels of historical dependence: Angel.R (no reference to previous cases), Angel.D (displays retrieved cases as context), and Angel.C (compares each retrieved case with the current query and returns an analysis as context). Prompts of this angels are provided in Appendix C.1. By aggregating independent diagnoses from these three agents and facilitating debate among their conclusions, our final diagnosis model, multi-Angels, bridges the gap between computational decision-making and the nuanced understanding essential for accurate psychiatric evaluations. The following parts introduce the main components of our agents:
Symptom Matching. To align client symptoms in medical records with DSM-5 diagnostic criteria, we process records and compute relevance between records and criteria using dense vector encoding  [^4] The processing steps of medical records are detailed in Appendix C.2.1, and the formulation of symptom matching is provided in Appendix C.2.2.. The BGE-M3 embedder [19] is employed for its strong semantic embedding capabilities. We retrieve the top-5 most similar criteria, returning their text, classification, and similarity scores. The tool does not diagnose but provides results for agent analysis, ensuring decisions integrate quantitative data and clinical expertise, mitigating over-reliance on single metrics.
@@
Scale Performance Analysis. We diagnose the presence of mood disorder using 16 key questions selected in Section 2.1 as the most mood-relevant items. Client performances are converted from numeric scores to textual descriptions based on question content and options. For agent interpretability, performances are reorganized into coherent descriptive paragraphs, enhancing analysis effectiveness (examples are provided in Appendix C.3).
Similar Cases Retrieval. To leverage clinical experience from similar cases, we develop two optional tools for retrieving medical records and scales with similar performance. After performing similarity retrieval, our tools generate different outputs tailored to the type of diagnosis agent in use. For Angel.R, this tool is intentionally excluded to minimize potential interference from the diagnostic outcomes of other cases. For Angel.D, the tool directly returns the retrieved cases for reference, enabling the agent to review and draw insights from them. For Angel.C, the tool conducts a detailed comparison of similarities and differences among the retrieved cases and returns an analysis text summarizing the findings. Consistent with the symptom matching tool, we employ BGE-M3 as the retriever. This approach ensures that the tool adapts to the specific needs of each diagnosis agent, enhancing the diagnostic process while maintaining flexibility and precision.
Multi-agent Diagnosis. To integrate insights from all Angels and improve diagnostics, Angel.R, Angel.D, and Angel.C first provide independent decisions and reasoning. A Judge Agent consolidates their inputs. If consensus is reached, the Judge outputs the diagnosis and reasoning. For disagreements, two Debate Agents are introduced: a Positive Agent, supporting a mood disorder diagnosis, and a Negative Agent, opposing it. Both Debate Agents and the Judge access symptom matching results, scale performances, relevant cases, and the Angels’ diagnosis and reasoning. In each debate round, the Positive Agent speaks first, followed by the Negative Agent. After each round, the Judge evaluates the arguments and decides whether to conclude the debate. If concluded, the Judge delivers the final diagnosis and supporting reasons. More details about the judge and a complete debate example are presented in Appendix C.4.1.
@@ 3 MoodSyn Dataset
The MoodSyn dataset addresses the critical need for clinically valid yet privacy-preserving data in computational psychiatry by introducing an open-source collection of 1,173 synthetic cases, as the statistics shown in Table 2. Each case captures the essential characteristics of psychiatric assessments through 25 carefully selected features, including 16 diagnostic questions, 8 standard scale scores, and expert-verified mood disorder labels. A data example is shown in Appendix D.1.
@@
Dataset Construction. The MoodSyn dataset is constructed on an advanced synthesis pipeline built upon the TabSyn framework [20], which integrates variational autoencoders with diffusion models through several technical innovations. The architecture employs adaptive tokenization and hierarchical encoding to maintain the complex relationships between clinical features, while a conditional denoising process preserves the nuanced symptom patterns characteristic of mood disorders. During generation, dynamic loss weighting ensures the mathematical consistency between individual question responses and their corresponding scale scores, and context-aware sampling maintains clinically meaningful feature correlations. The construction process is detailed in Appendix D.1.
Dataset Evaluation. Following Zhang et al. [20], we conduct a comprehensive evaluation of MoodSyn across five critical dimensions: statistical density, data quality, machine learning efficacy, privacy preservation, and logistic detectability (detailed in Appendix D.2). The results demonstrate exceptional fidelity to real clinical data, with MoodSyn accurately reproducing both core statistical patterns and complex symptom relationships while maintaining strong utility for machine learning applications. Specifically, our evaluation reveals high-density preservation of univariate and multivariate distributions matching the original data’s feature patterns, robust performance in downstream diagnostic prediction tasks comparable to real data, and near-indistinguishability in rigorous detection tests. Furthermore, MoodSyn provides stronger privacy guarantees than traditional anonymization approaches through its synthetic generation process. This unique combination of statistical fidelity, clinical validity, and privacy protection establishes MoodSyn as both a reliable analytical surrogate and a valuable resource for advancing AI applications in mental health research.
@@ 4.1 Experimental Settings
Baselines. To evaluate the effectiveness of our agent, we compare it against five baseline LLMs: LLaMA3-8B-Instruct [21], Mistral-7B-Instruct-v0.3 [22], GPT-4o (gpt-4o-2024-08-06) [23], DeepSeek-V3 [24], and medgemma-27b-text-it [25]. Each LLM is provided with a standardized input consisting of: (1) A combined medical record, presented as a unified narrative. (2) A summary of scale test results, listing the client’s scores across multiple psychological scales along with their clinical implications. The LLMs are prompted to determine whether the client has a mood disorder (including depression or bipolar disorder) and to provide a structured explanation in JSON format. An example of the prompt used for baseline models is shown in Figure 8.
Hyperparameters. The retrieval number $k$ for Angel.D and Angel.C is set to $k=5$ by default. For each model, we employ default parameter settings, utilizing official models for open-source LLMs obtained from Hugging Face or the API from the official website. These testing procedures take place on a computational infrastructure consisting of two NVIDIA A800 Tensor Core GPUs, equipped with 80GB of memory.
@@
More details about baseline prompts and the selection of retrieval number $k$ are provided in Appendix E.1.
@@ 4.2 Analysis of Experimental Results
The experimental results presented in Tables 3 and 4 demonstrate the clear advantages of this approach. We further observed that the generated reasoning demonstrates high quality, as detailed in Appendix E.2.1. All MoodAngels variants show substantial improvements over baseline LLMs, with even our foundational agent achieving 12.3% greater accuracy than GPT-4o, underscoring the effectiveness of our redesigned diagnostic process. A deeper examination of the agent comparisons reveals important insights about psychiatric diagnosis. The performance advantage of Angel.D over Angel.R confirms the clinical value of historical cases as reference points, while the slight dip observed with Angel.C serves as a caution against over-reliance on past cases. Particularly telling are those instances where only one agent succeeds in making the correct diagnosis while others fail, highlighting both the complexity of psychiatric assessment and the complementary nature of different diagnostic approaches.
We also conduct subgroup analysis on gender and age, confirming no significant bias in predictions (detailed in Appendix E.2.2). These findings collectively validate our multi-agent framework’s design, which synthesizes granular symptom analysis, structured diagnostic protocols, and balanced clinical experience integration through its collaborative debate mechanism. The framework’s ultimate performance superiority emerges from this sophisticated combination of innovations, each addressing specific limitations in conventional psychiatric assessment methods while working in concert to achieve more accurate and reliable diagnoses.
@@
Model Recall ACC MCC Macro F1
@@
$\text{Angel.R}_{GPT-4o}$ 0.840 0.920 0.829 0.913
$\text{Angel.D}_{GPT-4o}$ 0.845 0.923 0.837 0.917
$\text{Angel.C}_{GPT-4o}$ 0.848 0.914 0.814 0.906
$\text{multi-Angels}_{GPT-4o}$ 0.881 0.925 0.834 0.917
@@
$\text{Angel.R}_{Deepseek-V3}$ 0.863 0.927 0.841 0.920
$\text{Angel.D}_{Deepseek-V3}$ 0.864 0.920 0.823 0.911
$\text{Angel.C}_{Deepseek-V3}$ 0.858 0.922 0.829 0.914
$\text{multi-Angels}_{Deepseek-V3}$ 0.866 0.923 0.832 0.916
@@
Method Recall ACC MCC Macro F1
@@
$\text{Angel.R}_{Deepseek-V3}$ 0.778 0.8 0.601 0.798
$\text{Angel.D}_{Deepseek-V3}$ 0.787 0.807 0.615 0.805
$\text{Angel.C}_{Deepseek-V3}$ 0.816 0.821 0.642 0.821
$\text{multi-Angels}_{Deepseek-V3}$ 0.824 0.821 0.642 0.821